OpenStudy (anonymous):

6x – 9y = 16 2x – 3y = 7

OpenStudy (texaschic101):

elimination method.... 6x - 9y = 16 2x - 3y = 7 --->(-3) ---------------- 6x - 9y = 16 -6x + 9y = -21 (result of multiplying by -3) ----------------add 0 = - 5 (incorrect) There is NO SOLUTION to this system of equations